OpenStudy (anonymous):

3+12 -- divided by x 1 - 16 -- x^2

jhonyy9 (jhonyy9):

do you realy can understanding what you have wrote here ? please write it newly

OpenStudy (anonymous):

3 + 12/x = (3x + 12) / x 1 - 16/ x^2 = (x^2 - 16) / x^2 are the left hand expressions correct?

OpenStudy (anonymous):

no its not the question is 3 and then + 12/x divided by 1-16/x^2

jhonyy9 (jhonyy9):

12 16 3x+12 x*2 - 16 3x+12 x*2 3 + ---- / 1 - ---- =------- / --------- = ------ x ------- = x x*2 x x*2 x x*2 -16 x(3x+12) 3x(x+4) 3x =--------- = --------- = ----- x*2-16 (x-4)(x+4) x-4

OpenStudy (anonymous):

division now becomes (3x + 12) / x * x^2 / (x^2 - 16) = x * 3(x+4) / (x+4)(x-4) = 3x / (x-4)

OpenStudy (anonymous):

thank you so much guys i appreciate it =)